Answer
0.2
Work Step by Step
$x$ is the number chosen.
Since there are 10 numbers and each have an equal probability of being chosen, p = 0.1. Therefore, it follows:
p (X = 8 or X = 9) = 0.1 + 0.1 = 0.2
You need to log in to continue
You can help us out by revising, improving and updating this answer.
Update this answerAfter you claim an answer you’ll have 24 hours to send in a draft. An editor will review the submission and either publish your submission or provide feedback.